OpenStudy (anonymous):

Find the missing term of the geometric sequence. 1,250,_______, 50, ...

Parth (parthkohli):

The common ratio is 5..

Parth (parthkohli):

You multiply 50 with 50 to get the missing term.

Parth (parthkohli):

No, wait, 50 with 5.

Parth (parthkohli):

The common ratio is 5 then.

Parth (parthkohli):

50:x=x:1250 x^2 = 50 * 1250 x^2 = 62500 x = 250

Parth (parthkohli):

Get it?

Parth (parthkohli):

Common ratio means that(a common generalisation) \(\Large \color{purple}{\rightarrow a_{n -1}:a_{n}:: {a_n}:a_{n + 1} }\)

OpenStudy (anonymous):

Thanks

OpenStudy (campbell_st):

well you know the 3rd term and 1st term use thie to find the common ratio \[50 = 1250r^2\] \[r^2 = \frac{50}{1250}\] the common ratio is \[r = \frac{1}{5}\] then use \[T _{2}= 1250 \times (\frac{1}{5})^{2-1}\]

OpenStudy (anonymous):

the answer is 250

Parth (parthkohli):

Yes, yes, common ratio is actually 1/5... but the answer is correct.

OpenStudy (campbell_st):

lol... parth's solution needs a bit or work...

Parth (parthkohli):

Common ratio = Ratio of previous term to the current term

Parth (parthkohli):

@campbell_st ?

OpenStudy (campbell_st):

the common ratio is.... \[\frac{T _{2}}{T _{1}} = \frac{T _{3}}{T _{2}}=\frac{T _{4}}{T _{3}}...\]

OpenStudy (anonymous):

@campbell_st is his answer right?

OpenStudy (campbell_st):

the 2nd term is 250 1250 x 1/5

OpenStudy (anonymous):

that = 250